OpenStudy (kaylardz0405):

What is the value of the expression |a + b| + |c| when a = –5, b = 8, and c = –13? A. –26 B. 0 C. 16 D. –10

OpenStudy (kaylardz0405):

@mathmate

OpenStudy (kaylardz0405):

i had thought a but im not to sure

OpenStudy (mathmate):

hints: Step 1: substitute all the values of a, b and c into the expression. Step 2: Evaluate the expression within the | | (absolute value symbols). Step 3: Evaluate the absolute value functions, i.e. remove the | | symbols appropriately Step 4: Evaluate the resulting expression, that's your answer.

OpenStudy (kaylardz0405):

would the -5 turn into a postive or not

OpenStudy (mathmate):

For example, if a=4, b=6, c=-2 |a-b|+|c| =|4-6|+|-2|=|-2|+|-2|=2+2 = 4

OpenStudy (mathmate):

-5 does not turn into positive by itself. You need to complete step 2 first, i.e. reduce the a+b into a single value. Values turn positive only when there is a single number inside the | | symbols. See example above.

OpenStudy (kaylardz0405):

so then the answer is 16

OpenStudy (mathmate):

can you show your work so we can see how you worked it out?

OpenStudy (kaylardz0405):

yes

OpenStudy (kaylardz0405):

-5+8=3 13+3=16

OpenStudy (mathmate):

is c= 13 or -13?

OpenStudy (drigobri):

The expression is: |a + b| + |c| Change the values of a, b and c, the values they represent, in the case: a = –5, b = 8, and c = –13 That's: |-5 + 8| + |-13| The module (| x |) means that no matter the signal that the calculation result, it will always be positive. So: |-5 + 8| + |-13| |3| + |-13| 3 + 13 = 16 C = 16

OpenStudy (mathmate):

Sorry, c is -13, so 3+|c|=16, is correct.

OpenStudy (kaylardz0405):

thx